OpenStudy (anonymous):

Help please What is the sum of the first 5 terms of a geometric sequence with a1=10 and r=1/5? Express your answer as an improper fraction in lowest terms without using spaces

OpenStudy (anonymous):

a1=10. r=1/5 10 is starting term. 1/5 is the multiplier. 10 x (1/5) = 2 2 x (1/5) = 2/5 (2/5) x (1/5) = 2/25 (2/25) x (1/5) = 2/125 (2/125) x (1/5) = 2/625 add all those answers together using a common denominator of 625 to get the sum. Then simplify